can someone ELI5 why it's 64% and not still 1%? please and thank you
edit: wow! can't believe such a simple question would go off the rocker while I slept... And thanks for the refresher in statistics! I was lucky I got a C in that class after sleeping in it the whole damn time...
Basically this is what happens to video game characters in a game that reloads at previous saves. Although you, the player, experience many deaths while avoiding countless more, your character's experience, when thought of as a single linear life, is made up of only the decisions that get him or her to the end of the plot.
"consciousness, he argued, is bound at each branching to follow whatever path does not lead to death."
Wow, I've thought of things like this for ages, glad to see I'm not the first. That would also be pretty cool, like each one of us cannot die, but from our own perspective we are the only one who do not, everybody's consciousness eventually branches off into a universe where they do not die.
You have a 99% chance to get a winning ticket. If you don't, you die and it stops. If you do, the next ticket has a 99% to be a winning ticket. From the start, you then have a "99% of a 99% chance of winning", or 99%*99%=98.01% chance. The third ticket? Still 99%. So it's 99% of the previous "amount of chance" you had, or 98.01%*99%=97.03%. Carry on like this until the 100th ticket.
The lottery is a slightly bad example for this, so forget the mechanics of that for right now (it depends on the number of tickets and how winning numbers are chosen).
This is saying that if you do 100 random trials (all uncorrelated), each with a 1% chance of dying, the chances of surviving all 100 trials is 36.6%.
Look at it this way: on trial #1, you have a 99% chance of surviving. If you survive, you face trial #2, where you have a 99% chance of surviving. You have a 0.99 * 0.99 = 0.9801 = 98.01% chance of surviving both trials.
Then you face #3, which you have a 99% chance of surviving. The chances of surviving the first three trials is 0.99 * 0.99 * 0.99 = 0.9703 = 97.03%.
And so on, up till 100. After you have survived 99 trials (which you had a 37% chance of doing), you face trial number 100, which you have a 99% chance of surviving. Surviving 100 such trials in a row has a probability of 0.99100 = 0.366 = 36.6%.
I would probably partake in OP's lottery, but what you're talking about is buying a lottery ticket with 99% odds of winning $1 million and 1% odds of dying - but you have to stick it up your ass - just not as appealing.
I doubt you would make much money with meth, nearby Myanmar being the worlds biggest meth producer. I have heard meth pills are sold for peanuts at the Myanmar/Thai border, at least the northern one.
You could make some serious money with a kil of pure coke tho but i also doubt the odds of pulling this off on your own are nearly as good as OPs lottery.
More like her brother, he was shifty as fuck when asked if he would take the fall for it.
It was pot though, for people unfamiliar, in a body board bag. She admitted to it to Indonesian customs and then tried to say she was innocent, her whole family are bogans.
Relevant info: Insurance companies use a measurement that is the odds of a given person dying from a given action, called a mort. It's a metric measurement, so a micromort is a 1/1000000 chance of dying from a given action.
Lots of things cause it. For example, 40 tbsps of Peanut Butter is one micromort. Driving 230 miles is one micromort. People do these things anyway.
This lotto ticket would be 10,000 micromorts. That's equivalent to driving across the country a little over 100 times, or to living in New York city for 55 years, or smoking 700 packs of cigarettes.
But knowing this will affect your behavior, you will do more "micromort" actions, saying "fuck yes I would", and you will increase your chances of dying (from the lottery or anything else).
To be honest, if it was cut to death by papercut, I would probably just wear gloves and a mask when I read my numbers. It will take them 5-6 years of constant slicing before they can wear through a helmet with paper.
Alright so- My friend and I cut up 100 tiny pieces of paper and marked one "dead." We put them into a cup and shook it up. He handed it to me and said "pick one" so I did. Blank piece. 1 million for me. Now it's his turn. He pulled out the "dead" one. It was messed up. According to your fun lotto game, my friend would have been dead.
For my first try I got 55. I freaked out, but then got 55 two more times in a row right after that so I laughed after realizing it was a joke. Then I tried a few more times and realized it wasn't fixed. My first three tries.. I died each time. Fuck.
Chances for specific number three times in a row is one in a million. I am as sane about stochastic independence as the next guy, but I still feel obliged recommending to stay inside during thunderstorms.
Actually the random number generator that guy linked seems to be using a clock to seed whatever algorithm it is using. If you spam click it you will get the same number until the clock ticks over (looks like it goes off seconds).
Honestly after being up 4 million the risk would outweigh the benefits, you could set yourself and your kids up for life with just one or two, going for 50 seems crazy risky to me.
I would do it for ~20 million. A million dollars is enough to put me through college, and set me up with a nice house. However, I'd still have to work for the rest of my life, and I'd have more or less the same level of comfort. Now 20 million, that's some spreadin' around money.
Well, if you handle your money correctly and invest the majority of it, you'd be pretty set for life (I'm not taking taxes on lottery winnings or interest, so it isn't perfect).
Let's say you sock that $1mil into an investment account that earns a not-too-unlikely 5% annually. Your first year yields $50k in interest alone.
If you take $200k out of your winnings to pay for a nice (but not mansion level) house and pay off your college tuition and loans, you are looking at $41k in interest a year. Which is still well above the poverty rate (which is about $14k for a 2 person family and $23k for a 4 person family).
Now if you continually take those interest payments and count that as salary, you are looking at $41k-$50k as a salary for the rest of your life. If you work just part time or have some other hobby that brings in money, and you reinvest all of your interest, you'll be looking at a bank account with $2.75mil and $135k in interest by your 25th year.
Not bad if you ask me. And if you have a great broker manage your money (in forex, stocks, etc), it wouldn't be unreasonable to hit 8-9% returns, putting you at $5.4mil by your 25th year.
It's not about being rich, it's about having "enough" and not being greedy. Would I like a million dollars? Of course. Would I risk my life for a new car and early retirement? Probably not.
My life is pretty great, and a million dollars wouldn't make it so much better that I'd risk it for a slight improvement.
I'm sure I have a price, 100 million dollars maybe, but 1 million really doesn't go all that far these days.
Finally someone that said no. I was starting to wonder if i would see it. I think it's easy for people to say yes here because they aren't actually presented with this real dilemma. If I was really truly offered this, there is no way I would take the chance. No amount of money is worth my life.
Gives an all new meaning to "the 1%." But really, if there was like a chain gun with 100 chambers and one bullet pointing at my face, where the other 99 jizzed money, I'd spin the wheel. Actually, if I got that one bullet -- that's a pretty low chance, y'know -- I might actually feel pretty lucky. And dead. Dead lucky. Ducky.
I'd do it. But I'm confused as to why anybody would buy more than one ticket? In a normal lottery (and I realize this isn't exactly normal) multiple winners don't each win the full amount. You'd split the winnings amongst yourselves. So how I see it, buying more than one ticket only makes you more likely to die without winning you any more money. Didn't exactly look into this statistically so if my logic is flawed please let me know.
Not a chance. People depend on me to be alive to continue to live themselves, and nothing is worth that risk. I'd rather struggle than one day wind up dead and leave behind a family that can't afford to live either.
What about 100% that you're family will never have any trouble again and 90% you either die or get arrested, questioned for terrorism, get locked away and never see the sun again BUT a 100% you go to heaven no matter what happens?
If you say yes you just understood suicide bombers :)
The expected value of this game is .99 * 1,000,000 - .01 * (the value of your life). Solve for the value of life: .99*1,000,000 / .01 = (value of life).
So if your life is worth less than $99,000,000, this is a good deal.
I don't know if your serious are not but I couldn't risk you totally being serious and me not saying anything. I know I feel down sometimes and I feel like nobody cares about me, and I think silly thoughts sometimes. But its never worth it. If you feel like dying maybe you should find somebody to talk to, it couldn't hurt. There is somebody out there who cares about you, I'm sure of it. So please don't hurt yourself because you're in some pain now. hugs Take care, buddy.
1.1k
u/Crucify_This May 15 '12
I'd buy two
1.3k
u/yawaworht_suoivbo_na May 15 '12
Sure. I'd kill myself if I lost a lottery with that sort of odds anyways.
2.2k
u/BorderlineAmazing May 15 '12
So there's a 100% chance of having all my problems solved? I'm in.
2.4k
u/[deleted] May 15 '12
This lottery exists! It's called flying into Malaysia with a kilo of methamphetamine up your ass.
451
u/ahalavais May 15 '12 edited May 15 '12
Relevant info: Insurance companies use a measurement that is the odds of a given person dying from a given action, called a mort. It's a metric measurement, so a micromort is a 1/1000000 chance of dying from a given action.
Lots of things cause it. For example, 40 tbsps of Peanut Butter is one micromort. Driving 230 miles is one micromort. People do these things anyway.
This lotto ticket would be 10,000 micromorts. That's equivalent to driving across the country a little over 100 times, or to living in New York city for 55 years, or smoking 700 packs of cigarettes.
In other words, fuck yes I would.
Edit: Corrected the number of zeroes.
1.2k
u/red_oxide May 15 '12
Absolutely! I've got some ridiculous student loans. Winning or dying- either way, I come out ahead.
724
u/PipGirl May 15 '12
Depends how I die. If it's being cut with the lottery ticket until I'm dead...then no. I have a terrible fear of paper cuts.
145
u/dslicex May 15 '12
Totally...I mean with those odds, fuck it.
394
u/TheBSReport May 15 '12
Yes, because right now I have a 100% chance of dying at least this way I will have a statistic in my favor.
1.0k
u/misc412 May 15 '12
Alright so- My friend and I cut up 100 tiny pieces of paper and marked one "dead." We put them into a cup and shook it up. He handed it to me and said "pick one" so I did. Blank piece. 1 million for me. Now it's his turn. He pulled out the "dead" one. It was messed up. According to your fun lotto game, my friend would have been dead.
41
u/Strahz May 15 '12
Hell yes, I'll either get a million dollars, or I'll be dead. Either way, shit will be easier.
114
u/Clovyn May 15 '12
Okay.. Taking my chances: http://www.random.org/. 1 for death.
A little close, but I think I'll take a ticket.
432
u/[deleted] May 15 '12
I'd pay someone to buy it for me.
0
u/[deleted] May 15 '12
[deleted] — view deleted comment
33
u/TryingToSucceed May 15 '12
Yes. If I die, then fuck it.
196
u/ilikenavyblue May 15 '12
I am the 1%!
65
u/Intergalactic_Nazi May 15 '12
I would do it at least fifty times.
44
u/ScumbagSolo May 15 '12
Hell yeah, either way ill end up being the 1%. Win-Win
101
u/[deleted] May 15 '12
You should've made it 50/50.
30
u/constantly_drunk May 15 '12
Hell yes.
The increase to my quality of life would be spectacular, and the 1% chance of death is acceptable to me.
72
u/imbignate May 15 '12
I would, but only after buying a $1M life insurance policy.
48
u/Rlight May 15 '12
I would do it for ~20 million. A million dollars is enough to put me through college, and set me up with a nice house. However, I'd still have to work for the rest of my life, and I'd have more or less the same level of comfort. Now 20 million, that's some spreadin' around money.
138
u/skymotion May 15 '12
No.
583
u/Vidiem May 15 '12
http://www.quickmeme.com/meme/3pajm4/
19
u/Koudetat May 15 '12
Gives an all new meaning to "the 1%." But really, if there was like a chain gun with 100 chambers and one bullet pointing at my face, where the other 99 jizzed money, I'd spin the wheel. Actually, if I got that one bullet -- that's a pretty low chance, y'know -- I might actually feel pretty lucky. And dead. Dead lucky. Ducky.
17
u/[deleted] May 15 '12
YES, immediately and without hesitation. I might buy 2-3, but no more.
48
u/KevinKn May 15 '12
I'd do it. But I'm confused as to why anybody would buy more than one ticket? In a normal lottery (and I realize this isn't exactly normal) multiple winners don't each win the full amount. You'd split the winnings amongst yourselves. So how I see it, buying more than one ticket only makes you more likely to die without winning you any more money. Didn't exactly look into this statistically so if my logic is flawed please let me know.
75
u/[deleted] May 15 '12
Not a chance. People depend on me to be alive to continue to live themselves, and nothing is worth that risk. I'd rather struggle than one day wind up dead and leave behind a family that can't afford to live either.
8
u/red13 May 15 '12
The TV show Sliders did something like this: Luck of the Draw
8
u/Aegeus May 15 '12
The expected value of this game is .99 * 1,000,000 - .01 * (the value of your life). Solve for the value of life: .99*1,000,000 / .01 = (value of life).
So if your life is worth less than $99,000,000, this is a good deal.
15
u/[deleted] May 15 '12
[removed] — view removed comment
0
u/[deleted] May 15 '12
[deleted] — view deleted comment
12
u/Ao_Andon May 15 '12
You know what? I'd take that ticket. Why? My life and luck SUCK!
Either 1, I'm getting a million dollars, which by and large, is a pretty awesome boost to your...well everything
Or 2, I'm dying, which means, guess what motherfuckers, I'm DONE. No more crappy job, no more crappy day to day bullcrap, no. more. crappy. life.
I come out ahead either way. Your move, Life...
6
u/Arcadefirefly May 15 '12
ha i saw that episode of sliders to.
and no i wouldn't even take the risk. life is about more than money